OpenStudy (anonymous):

a tugboat goes 180miles upstream in 10 hours. the return trip downstream takes 5 hours. find the speed of the tugboat without a current and the speed of the current

OpenStudy (mathmate):

Upstream speed=180/10=18 mph Downstream speed=180/5=36 mps Boat speed = (36+18)/2=27 mph = (sum+difference)/2 Current speed=(36-18)/2=9 mph =(sum-difference)/2